OpenStudy (anonymous):

Solve for x: 5 over 2x + 15 over 2 = 15

OpenStudy (aroub):

5/2x +15/2 =15 5/2x+15=30 5/2x=15 x=3

OpenStudy (anonymous):

Thank you so much.:)